OpenStudy (anonymous):

A research study in 2000 found that there were 440 coyotes in a given region. The coyote population is expected to grow at a rate of 17% each year. In what year will the population first exceed 10,000?

OpenStudy (perl):

Let t = # of years since 2000 Then we have: $$ \Large { y = 440 ( 1 + .17)^t } $$

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you help me with one more problem as well

OpenStudy (anonymous):

Sodium-24 has a half-life of 15 hours. How many grams of a 900 g sample will remain after 90 hours? 28.125 g 14.0625 g 7.5 g 7.03125 g

OpenStudy (perl):

We can use the equation: $$ \LARGE { y = 900 \left( \frac12 \right)^\frac{t}{15} } $$